# Zero-Shot EEG-to-Gait Decoding via Phase-Aware Representation Learning

### v1

@@ 3.3.1 Session-Specific Prediction Heads
In the second stage, we retain only the pretrained EEG encoder and attach a prediction head to each source-domain session. For a given EEG embedding $\mathbf{z}_{e}$ belonging to session $s$, the corresponding session-specific head $h_{s}$ outputs a prediction $\hat{\mathbf{y}}=h_{s}(\mathbf{z}_{e})$ for the final motor frame (i.e., the gait pattern at the end of the 2-second window). Each session head is trained using the mean squared error (MSE) loss:
$\mathcal{L}_{\text{sup}}=\frac{1}{N}\sum_{i=1}^{N}\|\hat{\mathbf{y}}_{i}-\mathbf{y}_{i}\|^{2}$
@@ 3.3.2 Domain Weighting Layer and Mixture Prediction
To leverage inter-session relationships during inference, we introduce a domain weighting mechanism that adaptively fuses predictions from all source-domain heads except the one corresponding to the input session [yaoimproving]. Given an EEG embedding $\mathbf{z}_{e}\in\mathbb{R}^{d}$ from session $s$, the mixture prediction is computed as:
$\hat{\mathbf{y}}_{\text{mix}}=\sum_{k=1}^{N_{\text{src}}}\alpha_{k}(\mathbf{z}_{e})\cdot h_{k}(\mathbf{z}_{e})$
@@
$\boldsymbol{\alpha}=\text{softmax}\left(W_{a}\mathbf{z}_{e}+\mathbf{b}_{a}+\log(1-\mathbf{m})\right)$
Here, $W_{a}\in\mathbb{R}^{N_{\text{src}}\times d}$ and $\mathbf{b}_{a}\in\mathbb{R}^{N_{\text{src}}}$ are learnable parameters, and $\mathbf{m}\in\{0,1\}^{N_{\text{src}}}$ is a one-hot mask where $m_{s}=1$ denotes the current session. The additive masking term $\log(1-\mathbf{m})$ sets the attention logit of the input session to $\infty$- ∞, ensuring that $\alpha_{s}=0$. This excludes the intra-domain head from the mixture, promoting the learning of a cross-domain feature aggregation function.
@@ 3.3.3 Joint Optimization with $\mathcal{L}_{\text{sup}}$ and $\mathcal{L}_{\text{df}}$
We jointly train the session-specific heads and the domain weighting mechanism using two objectives: - $\mathcal{L}_{\text{sup}}$: ensures each session head captures reliable subject-specific gait mapping. - $\mathcal{L}_{\text{df}}$: ensures the mixture prediction $\hat{\mathbf{y}}_{\text{mix}}$ is close to ground truth:
$\mathcal{L}_{\text{df}}=\frac{1}{N}\sum_{i=1}^{N}\|\hat{\mathbf{y}}_{\text{mix},i}-\mathbf{y}_{i}\|^{2}$
The total loss of stage II is a weighted sum of supervised Loss and domain fusion loss:
$\mathcal{L}_{\text{stage II}}=\mathcal{L}_{\text{sup}}+\lambda_{\text{df}}\mathcal{L}_{\text{df}}$
We set $\lambda_{\text{df}}=1$ to equally balance the supervised objective and the domain fusion objective, encouraging the model to maintain high prediction accuracy within each session while simultaneously learning to generalize across domains through adaptive head mixing.
@@ 3.3.4 Inference via Head Mixture
$\hat{\mathbf{y}}_{\text{test}}=\sum_{i=1}^{N_{\text{src}}}\alpha_{i}\cdot h_{i}(\mathbf{z}_{e}^{tgt})$
where weights $\alpha_{i}$ are predicted by the domain weighting layer. This strategy enables zero-shot generalization by leveraging learned inter-session relationships without requiring any target-domain adaptation.
@@ 4.1.1 Gait-EEG Dataset
To investigate the brain mechanisms involved in walking, we collected a new dataset, Gait-EEG Dataset (GED) [fu2025eeg2gait], recording brain activity along with simultaneous lower-limb joint angles natural walk on level ground. The dataset contains the recordings from 50 able-bodied participants (25 males, 25 females; aged 21 to 46, mean age 28.4, standard deviation 5.2), with no history of neurological disorders or lower limb pathologies. Participants engaged in two independent level-ground walking experiment sessions, with every session comprising three identical walking blocks. Each block included approximately 40 trials, with each trial representing EEG signals and synchronized lower-limb joint angles as the participant walked straight for 7.7 meters. Sessions were spaced at least three days apart. The dataset includes synchronized recordings from a 60-channel active EEG, a 4-channel electrooculogram (EOG), along with measurements from six joint angle sensors (bilateral hips, knees, and ankles) [fu2025eeg2gait].
@@ 4.1.2 Open-access Dataset
To further validate our proposed method, we conducted additional experiments using the open-access Full Body Mobile (FBM) brain-body imaging dataset [brantley2018full]. The FBM dataset contains full-body motion capture data (66 markers) from approximately 10 walking trials performed by 10 able-bodied individuals under various gait conditions, including level ground, ramp, and stair walking. Data were collected using a wireless IMU-based motion capture system, enabling natural, unconstrained movement. To characterize lower-limb motor states, we selected the Z-axis joint angles of eight bilateral joints: hip, knee, ankle, and ball of the foot. A data segmentation stride of 50 ms was used.
@@ 4.3 Data Segregation
4.3.1 Gait-EEG Dataset
For NeuroDyGait, which is designed for cross-subject generalization, we split the 50 participants in the Gait-EEG dataset into training, validation, and test sets in an 8:1:1 ratio. Specifically, data from participants 1–40 were used for training, participants 41–45 for validation, and participants 46–50 for testing. All available data from each participant were included in their respective splits.
4.3.2 FBM Dataset
Similarly, in the FBM dataset, we used the complete data from participants 2–8 for training, participant 9 for validation, and participant 10 for testing. Data from participant 1 were excluded due to misalignment between the EEG and motion signals, which prevented reliable synchronization.
This subject-wise division ensures that the model is evaluated on entirely unseen individuals, emphasizing its generalization capability across subjects.
@@ 4.4 Evaluation Metric
Here, $y$ denotes the actual joint angle and $\hat{y}$ represents the predicted angle. The covariance between two variables $A$ and $B$ is denoted as $\operatorname{cov}(A,B)$, and $\sigma(A)$ is the standard deviation of $A$. $\bar{y}$ indicates the mean of the actual values. Each sequence represents data collected from a single trial over $n$ time steps, corresponding to $\frac{n}{100}$ seconds at a sampling rate of 100 Hz.
@@ 4.5 Implementation and Hyperparameter Settings
NeuroDyGait was implemented using the PyTorch library. Training was conducted using the Adam optimizer with default hyperparameter settings. A batch size of 512 was used, and training continued for a maximum of 50 epochs ($epoch_{\text{max}}$). Instead of using a fixed learning rate, we adopted a cosine learning rate scheduler with linear warm-up. The learning rate starts from zero and increases linearly during the initial warm-up phase, reaching the predefined maximum learning rate at the end of the warm-up period. After that, it gradually decreases to a minimum value following a cosine decay curve over the remaining training epochs. During the initial warm-up phase, spanning the first 2 epochs, the learning rate increases linearly from zero to the initial value. The total schedule is computed based on the number of training epochs and the number of iterations per epoch. The initial learning rate and minimum learning rate are listed in Table 1.
@@ 5 Results and Analysis
In this section, we conduct a comprehensive evaluation of the proposed framework across multiple key dimensions. We begin by assessing model performance, comparing our method against several state-of-the-art baselines on two benchmark datasets: GED [fu2025eeg2gait] and FBM [brantley2018full]. The results demonstrate consistent improvements in motion prediction accuracy, highlighting the effectiveness of our approach. We then examine cross-dataset transferability, evaluating the model’s generalization capability across different datasets, which underscores its robustness and adaptability in out-of-distribution settings. To further probe the model’s internal behavior, we visualize phase-specific EEG embeddings using t-SNE, revealing structured, phase-aligned representations that reflect the temporal specificity of the learned features. Additionally, we analyze the relationship between domain attention entropy and out-of-domain predictive error, illustrating how the domain-aware decoding mechanism leverages inter-session dynamics to enhance decoding performance. We also perform spatial analysis via saliency mapping to identify cortical regions critical to the model’s predictions. The results indicate that the model predominantly focuses on central sensorimotor channels, such as Cz and CP2, consistent with the neural correlates of lower-limb motor control. Finally, we present a series of ablation studies to systematically assess the contribution of key architectural and training components to the overall performance of the proposed framework.
@@ 5.1 Model Performance
In this section, we validate the performance of the proposed method on GED and FBM dataset and compared with several state-of-the-art deep learning and machine learning algorithms in the BCI domain, including ContraWR [yang2021self], FFCL [li2022motor], TSception [ding2022tsception], Temporal Convolutional Network (TCN) [ingolfsson2020eeg], ST-Transformer [song2021transformer], EEGConformer [song2023eeg], SPaRCNet [jing2023development], EEGNet [lawhern2018eegnet], and deepConvNet [schirrmeister2017deep] and EEG2GAIT [fu2025eeg2gait].
@@
The results of all metrics on the test sets are presented to evaluate the performance of NeuroDyGait in comparison to baseline methods on GED and FBM dataset. NeuroDyGait achieved an $r$ value of 0.673 ($R^{2}$ = 0.446, RMSE = 0.780) on GED dataset and an $r$ value of 0.259 ($R^{2}$ = 0.012, RMSE = 0.995) on FBM dataset, outperforming the baseline methods. All baseline methods were trained according to the strategy described in Section 4.3. A summary of the results for NeuroDyGait and the baseline methods on the different datasets is provided in Table 2 and 3. Across both GED and FBM datasets, NeuroDyGait consistently achieves either the best or second-best results on all evaluation metrics, demonstrating robust and competitive performance against existing state-of-the-art methods.
Table 2: The results of different methods on GED.
Method Pearson $r$uparrow $R^{2}$ Score $\uparrow$ RMSE $\downarrow$
ContraWR  [yang2021self] 0.2484 0.0515 1.0220
FFCL  [li2022motor] 0.3105 0.0743 1.0096
TSception  [ding2022tsception] 0.2783 0.0411 1.0271
EEGNet  [lawhern2018eegnet] 0.4450 0.1948 0.9413
TCN  [ingolfsson2020eeg] 0.2248 0.0448 1.0256
EEGConformer  [song2023eeg] 0.1279 -0.0099 1.0060
SPaRCNet  [jing2023development] 0.5946 0.3094 0.8713
ST-Transformer  [song2021transformer] 0.5764 0.3148 0.8679
deepConvNet  [schirrmeister2017deep] 0.6331 0.3928 0.8166
EEG2GAIT  [fu2025eeg2gait] 0.6348 0.3603 0.8385
NeuroDyGait 0.6725 0.4458 0.7799
↓: the lower the better; ↑: the higher the better.
The best results are highlighted in bold.
Table 3: The results of different methods on FBM.
Method Pearson $r$uparrow $R^{2}$ Score $\uparrow$ RMSE $\downarrow$
ContraWR  [yang2021self] 0.1315 -0.0926 1.0461
FFCL  [li2022motor] 0.1863 -0.0504 1.0259
TSception  [ding2022tsception] 0.1997 -0.0819 1.0409
EEGNet  [lawhern2018eegnet] $0$0.0622 1.0317
TCN  [ingolfsson2020eeg] 0.1348 -0.0304 1.0161
EEGConformer  [song2023eeg] 0.1279 -0.0099 1.0060
SPaRCNet  [jing2023development] 0.1347 -0.0669 1.0338
ST-Transformer  [song2021transformer] 0.1704 -0.0160 1.0111
deepConvNet  [schirrmeister2017deep] 0.2167 0.0063 0.9978
EEG2GAIT  [fu2025eeg2gait] 0.2189 0.0261 0.9879
NeuroDyGait 0.2585 0.0115 0.9952
↓: the lower the better; ↑: the higher the better.
The best results are highlighted in bold and the next best are marked using underlines.

### v2

@@ 3.3.1 Session-Specific Prediction Heads
In Stage II, we retain only the pretrained EEG encoder and attach a prediction head $h_{s}$ to each source-domain session. For a given EEG embedding $\mathbf{z}_{e}$ belonging to session $s$, the corresponding session-specific head $h_{s}$ outputs a prediction $\hat{\mathbf{y}}=h_{s}(\mathbf{z}_{e})$ for the final motor frame. Each session head is trained using the MSE loss:
$\mathcal{L}_{\text{sup}}=\frac{1}{N}\sum_{i=1}^{N}\|\hat{\mathbf{y}}_{i}-\mathbf{y}_{i}\|^{2}$ (10)
@@ 3.3.2 Domain Weighting Layer and Mixture Prediction
To leverage inter-session relationships during inference, we introduce a domain weighting mechanism that adaptively fuses predictions from all source-domain heads except the one corresponding to the input session [50]. Given an EEG embedding $\mathbf{z}_{e}\in\mathbb{R}^{d}$ from session $s$, the mixture prediction is computed as:
$\hat{\mathbf{y}}_{\text{mix}}=\sum_{k=1}^{N_{\text{src}}}\alpha_{k}(\mathbf{z}_{e})\cdot h_{k}(\mathbf{z}_{e})$ (11)
@@
$\boldsymbol{\alpha}=\text{softmax}\left(W_{a}\mathbf{z}_{e}+\mathbf{b}_{a}+\log(1-\mathbf{m})\right)$ (12)
Here, $W_{a}\in\mathbb{R}^{N_{\text{src}}\times d}$ and $\mathbf{b}_{a}\in\mathbb{R}^{N_{\text{src}}}$ are learnable parameters, and $\mathbf{m}\in\{0,1\}^{N_{\text{src}}}$ is a one-hot mask where $m_{s}=1$ denotes the current session. The additive masking term $\log(1-\mathbf{m})$ sets the attention logit of the input session to -$\infty$, ensuring that $\alpha_{s}=0$. This excludes the intra-domain head from the mixture, promoting the learning of a cross-domain feature aggregation function.
@@ 3.3.3 Joint Optimization with $\mathcal{L}_{\text{sup}}$ and $\mathcal{L}_{\text{df}}$
The session heads and domain weighting mechanism are jointly optimized with:
$\mathcal{L}_{\text{df}}=\frac{1}{N}\sum_{i=1}^{N}\|\hat{\mathbf{y}}_{\text{mix},i}-\mathbf{y}_{i}\|^{2}$ (13)
and the total Stage II loss:
$\mathcal{L}_{\text{stage II}}=\mathcal{L}_{\text{sup}}+\mathcal{L}_{\text{df}}$ (14)
@@ 3.3.4 Inference via Head Mixture
$\hat{\mathbf{y}}_{\text{test}}=\sum_{i=1}^{N_{\text{src}}}\alpha_{i}\cdot h_{i}(\mathbf{z}_{e}^{tgt})$ (15)
where weights $\alpha_{i}$ are predicted by the domain weighting layer. This strategy enables unseen-domain generalization by leveraging learned inter-session relationships without requiring any target-domain adaptation.
For clarity, Algorithm 1 provides the pseudocode of the complete two-stage training and inference process.
Input: EEG windows $X$, joint-angle labels $Y$, source domains $\{\mathcal{D}_{1},\ldots,\mathcal{D}_{N_{\text{src}}}\}$
Output: Predicted joint-angle trajectory $\hat{Y}$
Stage I: Phase-Aware Pretraining
foreach batch $(X,Y)$ do
 $Z\leftarrow f_{\theta}(X)$
 // EEG encoder
 $\hat{X}\leftarrow g_{\phi}(Z)$
 // Reconstruction
 $\hat{H}\leftarrow p_{\psi}(Z)$
 // Kinematic prediction
  Compute relative similarity scores $s(Z_{i},Z_{j})$
   Compute contrastive loss $\mathcal{L}_{\text{rcl}}$
  Compute total Stage I loss $\mathcal{L}_{\text{I}}=\|X-\hat{X}\|^{2}+\|H-\hat{H}\|^{2}+\mathcal{L}_{\text{rcl}}$
  Update $(\theta,\phi,\psi)$ via gradient descent
end foreach
Stage II: Domain Relation-Aware Decoding
Freeze encoder $f_{\theta}$
foreach source domain $\mathcal{D}_{k}$ do
   Train domain-specific decoder $d_{k}$
end foreach
Train attention module $A(\cdot)$ to output weights $\alpha=A(Z)$
Final prediction: $\hat{Y}=\sum_{k=1}^{N_{\text{src}}}\alpha_{k}\,d_{k}(Z)$
Inference
foreach incoming EEG window $X$ do
 $Z\leftarrow f_{\theta}(X)$
 $\alpha\leftarrow A(Z)$
 $\hat{Y}\leftarrow\sum_{k}\alpha_{k}\,d_{k}(Z)$
end foreach
return $\hat{Y}$
Algorithm 1 Two-Stage Training and Inference Pipeline of NeuroDyGait
@@ 4.1.1 Gait-EEG Dataset
To investigate the brain mechanisms involved in walking, we collected a new dataset, Gait-EEG Dataset (GED) [10], recording brain activity along with simultaneous lower-limb joint angles natural walk on level ground. The dataset contains the recordings from 50 able-bodied participants (25 males, 25 females; aged 21 to 46, mean age 28.4, standard deviation 5.2), with no history of neurological disorders or lower limb pathologies. Participants engaged in two independent level-ground walking experiment sessions, with every session comprising three identical walking blocks. Each block included approximately 40 trials, with each trial representing EEG signals and synchronized lower-limb joint angles as the participant walked straight for 7.7 meters. Sessions were spaced at least three days apart. The dataset includes synchronized recordings from a 60-channel active EEG, a 4-channel electrooculogram (EOG), along with measurements from six joint angle sensors (bilateral hips, knees, and ankles) [10].
@@ 4.1.2 Open-access Dataset
To further validate our proposed method, we conducted additional experiments using the open-access Full body Mobile brain-body imaging Dataset (FMD[4]. The FMD contains full-body motion capture data (66 markers) from approximately 10 walking trials performed by 10 able-bodied individuals under various gait conditions, including level ground, ramp, and stair walking. Data were collected using a wireless IMU-based motion capture system, enabling natural, unconstrained movement. To characterize lower-limb motor states, we selected the Z-axis joint angles of eight bilateral joints: hip, knee, ankle, and ball of the foot. A data segmentation stride of 50 ms was used.
@@ 4.3 Data Segregation
4.3.1 GED
For the Gait-EEG dataset, we adopted a 10-fold cross-subject evaluation to rigorously assess generalization. The 50 participants were evenly partitioned into ten disjoint folds, each containing data from five distinct subjects (e.g., subjects 1–5, 6–10, …, 46–50). In each round, one fold was used for testing, one for validation, and the remaining eight for training. This rotating scheme ensures that every subject serves as a test participant exactly once, providing a comprehensive measure of cross-subject robustness.
4.3.2 FMD
For the FMD, we followed a leave-one-subject-out (LOSO) protocol due to its smaller sample size of ten participants. In each iteration, data from nine subjects were used for training and one held-out subject for testing, while one additional subject was designated for validation within the training pool. Participant 1 was excluded from all folds because of EEG–motion desynchronization issues. This LOSO design enables an unbiased estimation of model generalization to completely unseen individuals.
These subject-wise cross-validation settings emphasize the model’s capacity to generalize across individuals rather than trials, aligning with the goal of robust and transferable gait decoding.
4.3.3 Segment Statistics
To provide a clearer understanding of the data volume used for model training and evaluation, we quantified the number of segmented EEG–gait windows generated per session after preprocessing and sliding-window segmentation.
For the GED dataset, each session contains on average 14,019.34 segments, with a standard deviation of 3,205.92 across all 100 sessions. For the FMD dataset, each session contains on average 32,877.44 segments, with a standard deviation of 5,047.82 across all 9 valid subjects.
These statistics provide a detailed view of the sample sizes available for the cross-subject evaluation protocols described above.
@@ 4.4 Evaluation Metric
Here, $y$ denotes the actual joint angle and $\hat{y}$ represents the predicted angle. The covariance between two variables $A$ and $B$ is denoted as $\operatorname{cov}(A,B)$, and $\sigma(A)$ is the standard deviation of $A$. $\bar{y}$ indicates the mean of the actual values. Each sequence represents data collected from a single trial over $n$ time steps, corresponding to $\frac{n}{{\color[rgb]{0,0,0}200}}$ seconds at a sampling rate of 200 Hz.
@@ 4.5 Implementation and Hyperparameter Settings
NeuroDyGait was implemented using the PyTorch library. Training was conducted using the Adam optimizer with default hyperparameter settings. A batch size of 512 was used, and training continued for a maximum of 50 epochs ($epoch_{\text{max}}$). Instead of using a fixed learning rate, we adopted a cosine learning rate scheduler with linear warm-up. The learning rate starts from zero and increases linearly during the initial warm-up phase, reaching the predefined maximum learning rate at the end of the warm-up period. After that, it gradually decreases to a minimum value following a cosine decay curve over the remaining training epochs. During the initial warm-up phase, spanning the first 2 epochs, the learning rate increases linearly from zero to the initial value. The total schedule is computed based on the number of training epochs and the number of iterations per epoch. The highest learning rate and minimum learning rate are listed in Table 2.
For clarity and reproducibility, all dataset-specific quantities referenced in the equations are explicitly summarized here. For both GED and FMD, we use a sequence length of $T=400$ samples (corresponding to 2 seconds at 200 Hz), $C=64$ EEG channels, and $J=6$ lower-limb joints. The number of source-domain sessions $N_{\text{src}}$ follows the leave-one-subject-out protocol: for GED, $N_{\text{src}}=49$; for FMD, $N_{\text{src}}=9$. All symbols and variable definitions are consolidated in Table 1 (Notation Table).
Stage I uses equal weighting for the reconstruction, prediction, and relative contrastive losses. As discussed in Section 6, jointly optimizing these objectives yields stable convergence without conflicting gradients. Sensitivity analysis further shows that varying the loss weights within moderate ranges results in less than a 2% change in decoding performance, indicating robustness to the choice of loss weights. Likewise, varying the embedding dimension $d$ between 64 and 256 also produced less than a 2% performance change, demonstrating low sensitivity to this hyperparameter.
@@ 5 Results and Analysis
In this section, we conduct a comprehensive evaluation of the proposed framework across multiple key dimensions. We begin by assessing model performance, comparing our method against several state-of-the-art baselines on two benchmark datasets: GED [10] and FMD [4]. The results demonstrate consistent improvements in motion prediction accuracy, highlighting the effectiveness of our approach. We then examine cross-dataset transferability, evaluating the model’s generalization capability across different datasets, which underscores its robustness and adaptability in out-of-distribution settings. To further probe the model’s internal behavior, we visualize phase-specific EEG embeddings using t-SNE, revealing structured, phase-aligned representations that reflect the temporal specificity of the learned features. Additionally, we analyze the relationship between domain attention entropy and out-of-domain predictive error, illustrating how the domain-aware decoding mechanism leverages inter-session dynamics to enhance decoding performance. We also perform spatial analysis via saliency mapping to identify cortical regions critical to the model’s predictions. The results indicate that the model predominantly focuses on central sensorimotor channels, such as Cz and CP2, consistent with the neural correlates of lower-limb motor control. Finally, we present a series of ablation studies to systematically assess the contribution of key architectural and training components to the overall performance of the proposed framework.
@@ 5.1 Model Performance
In this section, we validate the performance of NeuroDyGait on both the GED and FMD datasets and compare it against a comprehensive set of state-of-the-art deep learning and machine learning approaches for EEG-based decoding. The baseline models include ContraWR [49], FFCL [25], TSception [7], Temporal Convolutional Network (TCN) [18], ST-Transformer [40], EEGConformer [41], SPaRCNet [20], EEGNet [22], deepConvNet [37], and the recent state-of-the-art EEG2GAIT model [10].
We emphasize that EEG-to-gait regression is a highly specialized task, and only a limited number of domain-generalization techniques exist specifically for this setting. To ensure both fairness and relevance, we selected the strongest and most widely adopted EEG-based regression and representation-learning models currently available. In particular, the inclusion of EEGConformer (2023), SPaRCNet (2023), and EEG2GAIT (2025) ensures that our evaluation covers the most recent methodological advances in EEG decoding. These models collectively represent the current landscape of high-performing architectures for neural-based gait prediction, providing a rigorous benchmark for assessing the effectiveness of NeuroDyGait.
@@
The results of all evaluation metrics are reported to characterize NeuroDyGait’s performance relative to existing methods on both datasets. NeuroDyGait achieved an $r$ value of 0.6980 ($R^{2}$ = 0.4847, RMSE = 0.7329) on GED and an $r$ value of 0.2945 ($R^{2}$ = 0.0485, RMSE = 0.9743) on FMD, consistently outperforming all baseline approaches. All baseline models were trained following the strategy described in Section 4.3, ensuring a fair comparison.
A summary of quantitative performance across methods is provided in Table 3 and Table 4. Across both GED and FMD, NeuroDyGait achieves the best or second-best performance on all metrics, demonstrating robust generalization ability and substantial improvements over established EEG-based decoding architectures.
Table 3: Performance comparison of different methods on the GED dataset (mean $\pm$ std across folds).
Method Pearson $r$ $\uparrow$ $R^{2}$ Score $\uparrow$ RMSE $\downarrow$
ContraWR  [49] 0.3227 $\pm$ 0.1263 0.0830 $\pm$ 0.0834 0.9857 $\pm$ 0.0477
FFCL  [25] 0.4793 $\pm$ 0.1020 0.1946 $\pm$ 0.1133 0.9217 $\pm$ 0.0663
TSception  [7] 0.3664 $\pm$ 0.0793 0.1197 $\pm$ 0.0846 0.9653 $\pm$ 0.0471
EEGNet  [22] 0.4301 $\pm$ 0.1697 0.1917 $\pm$ 0.1476 0.9219 $\pm$ 0.0801
TCN  [18] 0.2608 $\pm$ 0.0522 0.0621 $\pm$ 0.0373 0.9977 $\pm$ 0.0205
EEGConformer  [41] 0.5292 $\pm$ 0.0817 0.2698 $\pm$ 0.1005 0.8774 $\pm$ 0.0623
SPaRCNet  [20] 0.6422 $\pm$ 0.0655 0.3756 $\pm$ 0.1037 0.8093 $\pm$ 0.0704
ST-Transformer  [40] 0.5963 $\pm$ 0.0795 0.3471 $\pm$ 0.1097 0.8282 $\pm$ 0.0710
deepConvNet  [37] 0.6904 $\pm$ 0.0679 0.4773 $\pm$ 0.0955 0.7392 $\pm$ 0.0717
EEG2GAIT  [10] 0.6962 $\pm$ 0.0764 0.4819 $\pm$ 0.1091 0.7387 $\pm$ 0.0969
NeuroDyGait 0.6980 $\pm$ 0.0742 0.4847 $\pm$ 0.1085 0.7329 $\pm$ 0.0809
↓: lower is better; ↑: higher is better. The best results are in bold; second best are underlined.
Table 4: Performance comparison of different methods on the FMD (mean $\pm$ std across subjects).
Method Pearson $r$ $\uparrow$ $R^{2}$ Score $\uparrow$ RMSE $\downarrow$
ContraWR  [49] 0.1838 $\pm$ 0.0490 0.0291 $\pm$ 0.0278 0.9846 $\pm$ 0.0152
EEGConformer  [41] 0.2443 $\pm$ 0.0842 -0.0514 $\pm$ 0.1132 1.0215 $\pm$ 0.0540
EEGNet  [22] 0.1827 $\pm$ 0.0986 0.0256 $\pm$ 0.0280 0.9861 $\pm$ 0.0153
FFCL  [25] 0.1545 $\pm$ 0.1122 0.0078 $\pm$ 0.0479 0.9947 $\pm$ 0.0233
EEG2GAIT  [10] 0.1984 $\pm$ 0.0966 0.0198 $\pm$ 0.0322 0.9886 $\pm$ 0.0174
SPaRCNet  [20] 0.1811 $\pm$ 0.0349 -0.0379 $\pm$ 0.0288 1.0175 $\pm$ 0.0145
ST-Transformer  [40] 0.1716 $\pm$ 0.0258 -0.0367 $\pm$ 0.0281 1.0172 $\pm$ 0.0134
TCN  [18] 0.0709 $\pm$ 0.0312 -0.0136 $\pm$ 0.0151 1.0061 $\pm$ 0.0086
TSception  [7] 0.1698 $\pm$ 0.0512 -0.3285 $\pm$ 0.6594 1.1229 $\pm$ 0.2229
deepConvNet  [37] 0.2185 $\pm$ 0.0881 -0.0313 $\pm$ 0.1034 1.0124 $\pm$ 0.0505
NeuroDyGait 0.2945 $\pm$ 0.1131 0.0485 $\pm$ 0.0342 0.9743 $\pm$ 0.0185
↓: lower is better; ↑: higher is better. The best results are in bold; second best are underlined.